a 1.2 x10^3 kg block starting from rest is accelerated by a 2.20x10^3 N force for a distance of 50.0 m along a horizontal surface
1
Expert's answer
2020-12-15T11:47:25-0500

m = 1200kg - mass of block

F = 2200N - force

S = 50m - distance

a block is accelerated by force 2200N;

A - is a work under block "A = F * S =" "2200N * 50m = 110000N*m = 110kJ;"

a - accelerating "a = \frac{F}{m} = \frac{2200N}{1200kg} = 1.83 \frac{m}{s^2}"